OpenStudy (anonymous):

If f(x) = x + 2, find f(3).

OpenStudy (anonymous):

f(x) = (x) + 3. f(_) = (_) +3 whatever is in the parentheses on the left goes in the parentheses on the right. that's function notation.

OpenStudy (zzr0ck3r):

put 3 where you see x

OpenStudy (anonymous):

thx zz

OpenStudy (anonymous):

???

OpenStudy (anonymous):

f(x) = x + 2 f(3) = 3 + 2 f(3) = ______

OpenStudy (anonymous):

@pgpilot326 where did you get f(x) = x + 3 ??

OpenStudy (anonymous):

sorry... typo. should have been a 2

OpenStudy (anonymous):

haha yeah.